# Signal-SGN: A Spiking Graph Convolutional Network for Skeleton Action Recognition via Learning Temporal-Frequency Dynamics

### v3

@@ V Conclusion
Appendix
In the appendix, we first provide the supplemental detail description of Legendre Polynomial-Based Filters and pseudocode of the proposed methods, followed by a detailed introduction to the dataset. Then, we describe the hyperparameter settings and training details.
V-A Legendre Polynomial-Based Filters
@@
To apply these polynomials to time-domain signals sampled over $[0,T]$, we first normalize the time index:
$u=\frac{t}{T},\quad u\in[0,1].$ (24)
@@
$\Lambda_{0}[k,t]=\sqrt{2k+1}\,P_{k}(2u-1),$ (26)
$\Lambda_{1}[k,t]=\sqrt{2k+1}\,P_{k}(2(u+0.5)-1)=\sqrt{2k+1}\,P_{k}(2u).$ (27)
@@
$\Gamma_{0}[k,t]=\sqrt{2}\,\Lambda_{0}[k,t],$ (28)
$\Gamma_{1}[k,t]=\sqrt{2}\,\Lambda_{1}[k,t].$ (29)
Taken together, these four filters $(\Lambda_{0},\Lambda_{1},\Gamma_{0},\Gamma_{1})$ can decompose the signal into distinct frequency bands, enabling a nuanced analysis of spike-like features. By carefully mapping temporal samples into the Legendre polynomial domain, employing a normalized and shifted set of low-pass filters, and incorporating a high-pass counterpart, this approach provides a stable, orthogonal, and frequency-selective analytical framework. This framework facilitates effective characterization, improved interpretability, and subsequent analyses of time-varying signals in a multi-resolution context.
V-B Algorithm
@@
V-B1 Spiking Neuron Dynamics Simulation
@@
$\tau\frac{dV(t)}{dt}=-\left(V(t)-V_{\text{rest}}\right)+R\cdot G_{i}(t),$ (30)
@@ V-B2 Iterative Decomposition
1:  $X_{0}\leftarrow X_{\text{extended}}$
@@
3:     $X_{\text{even}}^{(j)}\leftarrow X_{j}[::2,:,:]$
4:     $X_{\text{odd}}^{(j)}\leftarrow X_{j}[1::2,:,:]$
@@
The iterative decomposition process begins with the extended spike-form feature tensor $X_{\text{extended}}$, which serves as the initial input. The tensor is initialized as $X_{0}$, providing the starting representation for decomposition across $J$ levels.
@@
$D_{j}=\Gamma_{0}\cdot X_{\text{even}}^{(j)}+\Gamma_{1}\cdot X_{\text{odd}}^{(j)},$ (31)
$S_{j}=\Lambda_{0}\cdot X_{\text{even}}^{(j)}+\Lambda_{1}\cdot X_{\text{odd}}^{(j)}.$ (32)
@@
Upon completion, the algorithm outputs a set of detail coefficients $\{D_{j}\}_{j=0}^{J-1}$ and the final scaling coefficients $S_{J}$, collectively encapsulating the hierarchical temporal structure of the input. The complete iterative decomposition process is outlined in Algorithm 2, providing a systematic and computationally efficient method for processing spike-form freature.
V-C Datasets
V-C1 NTU-RGB+D
NTU-RGB+D[30] is a comprehensive dataset widely used in the field of skeleton-based action recognition. It contains 60 action classes and 56,000 video clips, making it a robust benchmark for evaluating action recognition models. The dataset provides detailed 3D skeleton data, RGB frames, depth maps, and infrared sequences, covering a diverse range of daily activities such as eating, walking, and interacting with objects. The skeleton data consists of 25 joints per subject, captured using the Microsoft Kinect v2 sensor, which allows for precise modeling of human movement. This dataset is particularly valuable for researchers because it includes a variety of subjects performing actions from multiple viewpoints, offering a rich source of information to develop and test models capable of understanding complex human behaviors.
V-C2 NTU-RGB+D 120
NTU-RGB+D 120[31] is an extended version of the original NTU-RGB+D dataset, expanding the number of action classes from 60 to 120 and increasing the total number of video clips to over 114,000. This extension includes additional action categories that encompass more complex and subtle human interactions, such as ”taking a selfie” and ”writing on a board.” With 106 distinct subjects and more diverse camera angles, NTU-RGB+D 120 offers a more challenging and comprehensive benchmark for evaluating the scalability and generalization capabilities of action recognition models. The dataset retains the detailed skeleton tracking of its predecessor, providing 25 joints per person and ensuring that the data remains highly informative for developing models that require an in-depth understanding of spatial-temporal dynamics in human actions. As one of the largest skeleton-based action recognition datasets available, NTU-RGB+D 120 is instrumental in pushing the boundaries of research in this field.
V-C3 NW-UCLA
The NW-UCLA dataset[32] is a smaller but uniquely challenging dataset consisting of 10 action classes and 1,494 video clips. It is specifically designed to test the robustness of action recognition models against viewpoint variations. The dataset features actions like jumping, waving, and running, captured from multiple viewpoints using three Kinect cameras placed at different angles. Each action is performed by 10 subjects, offering diverse skeleton data with 20 joints per person. This setup mimics real-world scenarios where actions are observed from different perspectives, making it a critical dataset for evaluating how well models can maintain performance when faced with varying observational conditions. The NW-UCLA dataset is ideal for researchers aiming to enhance the adaptability and resilience of their models to changes in camera viewpoints, which is essential for applications in surveillance, human-computer interaction, and robotics.
V-D Model Implementation and Configuration
The models used for conducting experiments are implemented using PyTorch [^1] https://github.com/rwightman/pytorch-image-models, a classical deep learning framework known for its flexibility and efficiency, and SpikingJelly [^2] https://github.com/fangwei123456/spikingjelly, a specialized library for spiking neural network research. Our implementation leverages the powerful features of these tools to effectively model and simulate the complex dynamics of spiking neural networks. The training process is conducted on four NVIDIA V100 GPUs, which provide the necessary computational power and memory to handle the large datasets and intricate computations involved in our experiments. This setup ensures that our models are trained efficiently, allowing us to explore the full potential of our approach in a scalable and robust manner.
TABLE IV: The parameter Settings of the Backbone
Level of Layers Channel Size
Input Output
1 3 64
2 64 64
3 64 128
4 128 256
In this section, we also give the specific hyperparameters of the model and training settings in all experiments. Table IV. presents the parameter settings of the backbone network, detailing the input and output channel sizes across different layers to illustrate the model’s architectural configuration.
The various hyperparameter settings used in our experiments are also shown in the TableV.
TABLE V: Hyper-parameter training settings of Signal-SGN.
Parameter NTU-RGB+D NTU-RGB+D NW-UCLA
60 120
Learning Rate 0.1/0.1 0.1/0.1 0.1/0.1
Decay rate 0.1/0.1 0.1/0.1 0.1/0.1
Batch Size 64 64 128
Training Epochs 90 90 90
Step 50 50 50
Dropout Rate 0.3 0.3 0.3
Weight Decay 1e-4 1e-4 1e-4
$\omega$ 0.03 0.03 0.03
Optimizer SGD SGD SGD
V-E Power Consumption
We assume that multiply-and-accumulate (MAC) and accumulate (AC) operations are implemented on the 45 nm technology node with $E_{MAC}$ = 4.6 pJ and $E_{AC}$ = 0.9 pJ. The $\mathrm{FLOPs}$ for Signal-SGN and all reproduced models are calculated using FVCore [^3] https://github.com/facebookresearch/fvcore, providing accurate estimates for each model component. These FLOPs are then combined with the firing rate and spike time steps to compute SOPs as:
$\mathrm{SOPs}(l)=fr\times T\times\mathrm{FLOPs}(l),$ (33)
The model’s theoretical energy consumption is described as:
$\begin{gathered}E_{\text{Signal-SGN }}=E_{EAC}\times\mathrm{FL}_{\mathrm{SNN}\text{ Conv }}^{1}+\\ E_{AC}\times\left[\left(\sum_{n=2}^{N}\mathrm{SOP}_{\mathrm{SNN}\text{ Conv }}^{n}+\sum_{m=1}^{M}\mathrm{SOP}_{\mathrm{SNN}\text{ FC }}^{m}\right.\right.\\ \left.\left.+\sum_{l=1}^{L}\mathrm{SOP}_{\mathrm{SNN}\text{ SSA }}^{l}+\sum_{f=1}^{F}\mathrm{SOP}_{\mathrm{SNN}\text{ FFT }}^{f}\right)+\mathrm{SOP}_{\mathrm{SNN}\text{ MWTF }}\right]\end{gathered}$ (34)
where $\mathrm{FL}_{\mathrm{SNN}\text{ Conv }}^{1}$ is the first layer to encode skeleton into spike-form. Then the SOPs of $N$ SNN Conv layers, $M$ SNN Fully Connected Layer (FC), $L$ SSA , $F$ SSA and MWTF are added together are added together and multiplied by $E_{AC}$. The theoretical energy consumption for ANN and SNN models is calculated as follows:
For ANNs, the energy consumption of a block $b$ is given by:
$\text{Power}(b)=4.6\,\mathrm{pJ}\times\mathrm{FLOPs}(b).$ (35)
For SNNs, the energy consumption of a block $b$ is given by:
$\text{Power}(b)=0.9\,\mathrm{pJ}\times\mathrm{SOPs}(b).$ (36)
V-F More detail of Ablation Studies
In this section, we provide additional details regarding the ablation studies conducted in our research. These experiments are designed to investigate the contribution of various components within our framework and to demonstrate the effectiveness of specific design choices. By systematically removing or altering key elements, we aim to provide a deeper understanding of their impact on the overall performance.
3D coordinate normalization
3D skeleton data often spans multiple quadrants and includes negative values, as shown in Figure 2, creating challenges for SNNs, which are inherently designed for non-negative numerical inputs. Negative coordinates lead to significant information loss and degrade model performance. To address this limitation, 3D coordinate normalization is applied to map skeleton data into a fixed and bounded range, ensuring numerical consistency and preserving essential information. Normalization facilitates stable data propagation through the network by mitigating discrepancies arising from varying magnitudes and signs.
Figure 2: 3D Skeleton Visualization
The analysis focuses primarily on accuracy as the evaluation metric, as changes in parameters and FLOPs introduced by normalization are minimal. By isolating the impact of 3D coordinate normalization on accuracy, the study demonstrates that normalization significantly enhances model performance without introducing substantial computational overhead. Controlled experiments under various module configurations reveal consistent accuracy improvements, underscoring this approach’s effectiveness in processing complex skeleton data. Detailed results presented in the accompanying table highlight the practical benefits of normalization in achieving higher accuracy with negligible computational trade-offs.
TABLE VI: Impact of 3D Coordinate Normalization on Accuracy
3D Coord. Norm. 1D-SGC FSC MWTF Acc (%)
- - - 69.7
- - 77.0
- 79.3
- - 70.5
- 78.3
80.5
Figure 3: Comparison of confusion matrices for full-module integration setup with and without 3D coordinate normalization
The results in Table 3 demonstrate the significant role of 3D coordinate normalization in improving accuracy across all module configurations. For the baseline configuration (1D-SGC only), normalization increases accuracy by 0.8%, while in more complex setups with FSC and MWTF modules, the improvements reach 1.3% and 1.2%, respectively. This indicates that normalization effectively stabilizes the input data by addressing inconsistencies in coordinate ranges, enabling the model to better utilize its capacity. Importantly, the improvements in accuracy are achieved without notable changes in parameters or computational cost, highlighting normalization as a practical and efficient enhancement for processing 3D skeleton data. To further analyze the impact of 3D coordinate normalization, we selected the full-module integration setup (MWTF enabled) from the third and sixth rows of Table VI and visualized the corresponding confusion matrices.
Different Settings for MWTF
To investigate the impact of different configurations of $k$ (frequency resolution) and $J$ (decomposition levels) on the performance of the MWTF module within Signal-SGN, we conduct a series of controlled experiments. The overall architecture of Signal-SGN is kept unchanged, ensuring that any observed variations in performance can be attributed solely to changes in the MWTF parameters. By systematically adjusting $k$ and $J$, we aim to analyze how the trade-off between frequency granularity and decomposition depth affects the network’s ability to extract and fuse temporal features effectively.
TABLE VII: Comparison of experimental results under different MWTF settings ($k$ and $J$)
MWTF $k$ $J$ Param. (M) FLOPs/SOPs (G) Acc (%)
- - - 1.65 1.60/0.312 78.3
4 1 1.68 1.61/0.313 78.9
8 1 1.73 1.62/0.314 79.8
8 2 1.74 1.62/0.315 80.3
8 3 1.74 1.62/0.317 80.5
16 3 1.83 1.64/0.320 80.5
The experimental results in Table VII highlight the impact of varying $k$ (frequency resolution) and $J$ (decomposition levels) on the performance of the MWTF module within Signal-SGN. With the overall architecture unchanged, the table isolates the influence of MWTF parameters, providing insight into their effects on accuracy, computational complexity, and parameter efficiency.
When MWTF is disabled, the model achieves a baseline accuracy of 78.3% with a parameter count of 1.65M and computational costs of 1.60/0.312 G (FLOPs/SOPs). Enabling MWTF with different $k$ and $J$ configurations demonstrates significant improvements. As $k$ increases from $4$ to $8$, accuracy improves substantially, with the model achieving 79.8% ($k=8,J=1$). Further increasing $k$ to $16$ yields diminishing returns, with accuracy peaking at 80.5% for $k=16,J=3$.
Similarly, increasing $J$ allows for deeper decomposition of temporal features, which enhances performance up to a point. For example, with $k=8$, accuracy improves from 79.8% ($J=1$) to 80.7% ($J=3$). This configuration achieves the best accuracy while maintaining a parameter count of 1.74M and computational cost of 1.62/0.317 G. Further increasing $k$ to $16$ ($J=3$) slightly reduces accuracy to 80.5%, indicating that excessively high frequency resolution or decomposition depth might lead to redundant features or overfitting.
Overall, the results demonstrate that the MWTF module significantly enhances the performance of Signal-SGN when configured optimally. The configuration $k=8,J=3$ emerges as the most effective setting, balancing frequency granularity, decomposition depth, and computational efficiency, and achieving the best accuracy of 80.7%. This underscores the importance of selecting appropriate MWTF parameters to maximize feature extraction and model performance.

### v5

@@ 5. Conclusion
Appendix A Power Consumption
We assume that multiply-and-accumulate (MAC) and accumulate (AC) operations are implemented on the 45 nm technology node with $E_{\text{MAC}}$ = 4.6 pJ and $E_{\text{AC}}$ = 0.9 pJ (Horowitz, 2014). The $\mathrm{FLOPs}$ for Signal-SGN and all reproduced models are calculated using FVCore [^3] https://github.com/facebookresearch/fvcore, providing accurate estimates for each model component. These FLOPs are then combined with the firing rate and spike time steps to compute SOPs as:
(18) $\mathrm{SOPs}(l)=fr\times T\times\mathrm{FLOPs}(l),$
The model’s theoretical energy consumption is described as, according to (Tavanaei et al., 2019; Kundu et al., 2021; Zheng et al., 2021; Panda et al., 2020; Yao et al., 2021; Zhou et al., 2022):
(19) $\begin{gathered}E_{\text{Signal-SGN }}=E_{\text{EAC}}\times\mathrm{FL}_{\mathrm{SNN}\text{ Conv }}^{i}+\\ E_{\text{AC}}\times\left[\left(\sum_{n={\hat{n}}}^{N}\mathrm{SOP}_{\mathrm{SNN}\text{ Conv }}^{n}+\sum_{m=1}^{M}\mathrm{SOP}_{\mathrm{SNN}\text{ FC }}^{m}\right.\right.\\ \left.\left.+\sum_{l=1}^{L}\mathrm{SOP}_{\mathrm{SNN}\text{ SSA }}^{l}+\sum_{f=1}^{F}\mathrm{SOP}_{\mathrm{SNN}\text{ FFT }}^{f}\right)+\mathrm{SOP}_{\mathrm{SNN}\text{ MWTF }}\right]\end{gathered}$
where $\mathrm{FL}_{\mathrm{SNN}\text{ Conv }}$ is the first layer to encode skeleton into spike-form. Then the SOPs of $N$ SNN Conv layers, $M$ SNN Fully Connected Layer (FC), $L$ SSA layer , $F$ FFT layer and MWTF are added together are added together and multiplied by $E_{\text{AC}}$. The theoretical energy consumption for ANN and SNN models is calculated as follows:
For ANNs, the energy consumption of a block $b$ is given by:
(20) $\text{Power}(b)=E_{\text{EAC}}\times\mathrm{FLOPs}(b).$
For SNNs, the energy consumption of a block $b$ is given by:
(21) $\text{Power}(b)=E_{\text{AC}}\times\mathrm{SOPs}(b).$
Appendix B Legendre Polynomial-Based Filters
@@
To apply these polynomials to temporal-domain signals sampled over $[0,T]$, we first normalize the time index:
(24) $u=\frac{t}{T},\quad u\in[0,1].$
@@
(26) $\Lambda_{0}[k,t]=\sqrt{2k+1}\,P_{k}(2u-1),$
(27) $\Lambda_{1}[k,t]=\sqrt{2k+1}\,P_{k}(2(u+0.5)-1)=\sqrt{2k+1}\,P_{k}(2u).$
@@
(28) $\Gamma_{0}[k,t]=\sqrt{2}\,\Lambda_{0}[k,t],$
(29) $\Gamma_{1}[k,t]=\sqrt{2}\,\Lambda_{1}[k,t].$
Taken together, these four filters $(\Lambda_{0},\Lambda_{1},\Gamma_{0},\Gamma_{1})$ can decompose the signal into distinct frequency bands, enabling a nuanced analysis of spike-like features. By carefully mapping temporal samples into the Legendre polynomial domain, employing a normalized and shifted set of low-pass filters, and incorporating a high-pass counterpart, this approach provides a stable, orthogonal, and frequency-selective analytical framework. This framework facilitates effective characterization, improved interpretability, and subsequent analyses of temporal-varying signals in a multi-resolution context.
Appendix C Algorithm
@@
C.1. Spiking Neuron Dynamics Simulation
@@
(30) $\tau\frac{dV(t)}{dt}=-\left(V(t)-V_{\text{rest}}\right)+R\cdot G_{i}(t),$
@@ C.2. Iterative Decomposition
1:  $\widetilde{X}_{0}\leftarrow\widetilde{X}$
@@
3:     $X_{\text{even}}^{(j)}\leftarrow\widetilde{X}_{j}[::2,:,:]$
4:     $X_{\text{odd}}^{(j)}\leftarrow\widetilde{X}_{j}[1::2,:,:]$
@@
The iterative decomposition process begins with the extended spike-form feature tensor $\widetilde{X}$, which serves as the initial input. The tensor is initialized as $X_{0}$, providing the starting representation for decomposition across $J$ levels.
@@
(31) $D_{j}=\Gamma_{0}\cdot X_{\text{even}}^{(j)}+\Gamma_{1}\cdot X_{\text{odd}}^{(j)},$
(32) $S_{j}=\Lambda_{0}\cdot X_{\text{even}}^{(j)}+\Lambda_{1}\cdot X_{\text{odd}}^{(j)}.$
@@
Upon completion, the algorithm outputs a set of detail coefficients $\{D_{j}\}_{j=0}^{J-1}$ and the final scaling coefficients $S_{J}$, collectively encapsulating the hierarchical temporal structure of the input. The complete iterative decomposition process is outlined in Algorithm 2, providing a systematic and computationally efficient method for processing spike-form feature.
Appendix D Supplementary Analysis
D.1. Analysis of Window function
(a) Window weights at initialization.
(b) Window weights after training.
Figure 8. Visualization of the learned joint-wise window functions $w_{V}(v)$ across FSC layers.
To analyze the role of the learnable joint-wise window function $w_{V}(v)$, we visualize its values across all FSC layers at both the beginning and end of training, as shown in Figure 8. Initially, all layers exhibit near-zero or uniform weights, especially in deeper layers like Layer 4, where the window is almost flat and close to zero, as shown in Figure 8(a). This indicates that the network starts without any strong spatial preference in the frequency modulation stage.
After training, however, the learned windows evolve into highly diverse and non-uniform patterns across joints,as shown in Figure 8(b). The early layers (e.g., Layer 1) show sharp variations, possibly focusing on localized joints, while deeper layers develop smoother yet still distinctive profiles. This progression suggests that learning a fully adaptive window function allows the network to emphasize or suppress specific joints based on spatiotemporal dynamics, with joint-wise spectral weighting acting as a spatial prior to guide frequency decomposition according to human motion structure and semantics.
D.2. Analysis of Learnable Fusion Parameter $\beta$
Figure 9. Joint evolution of training loss and the learnable fusion weight $\beta$.
To complement the ablation study on the MWTF module, we provide an analysis of the learnable fusion weight $\beta$ in the final classification stage, as defined in Equation (17). Figure 9 illustrates the joint evolution of the training loss and the value of $\beta$ throughout training. We observe that $\beta$ fluctuates significantly during the initial 20 epochs, reflecting the model’s attempt to dynamically balance the temporal backbone output and the spectral feature $\hat{X}$ extracted by MWTF. As the training loss steadily declines, $\beta$ gradually stabilizes. By the end of training, $\beta$ converges to approximately 0.42, indicating that the model assigns moderate but consistent weight to the spectral branch. This behavior suggests that while the primary classification signal originates from the SN-processed temporal pathway, the MWTF-derived spectral features provide substantial complementary information. The learnable fusion mechanism, governed by $\beta$, thus plays a key role in adaptively integrating multiscale dynamics for robust action recognition.
D.3. Frequency-Domain Analysis of MWTF
To further interpret the effectiveness of the proposed spiking cross-attention fusion across wavelet decomposition levels, we perform a joint-wise spectral analysis using the example of the ”drinking water” action on NTU-RGB+D. As shown in Figure 10, each subfigure presents the magnitude spectrogram obtained from the FFT of a single channel with the highest temporal variance.
The final cross-attentively fused output $\hat{X}$ in Figure 10(a) exhibits an enriched spectrum that combines both low- and high-frequency components. Compared to the low-frequency-dominated baseline $X^{(L)}$ in Figure 10(b), $\hat{X}$ demonstrates enhanced high-frequency responses particularly at joints 6 (right shoulder), 8 (right elbow), 10 (right wrist), and 12 (right hand), which are anatomically responsible for raising the arm and bringing the hand to the mouth during drinking.
Figures 10(c)–(f) illustrate the spiking cross-attentively fused detail coefficients $\hat{D}_{0}$ to $\hat{D}_{3}$, representing progressively higher frequency bands as the decomposition level increases. Specifically, $\hat{D}_{0}$ and $\hat{D}_{1}$ capture mid-frequency activations in joints 6, 8, and 10, reflecting shoulder and elbow articulation during arm lifting. The component $\hat{D}_{2}$ highlights more localized and rapid changes in joints 10 and 12, aligning with the wrist-to-mouth trajectory. As the final decomposition level, $\hat{D}_{3}$ is generated via upsampling of the residual high-frequency component and retains only coarse responses—mainly from trunk-related joints (e.g., joints 2 and 3)—representing global pose adjustments.
Figures 10(g)–(i) show the corresponding scaling coefficients $\hat{S}_{0}$ to $\hat{S}_{2}$, also fused through spiking cross-attention (cf. Eq. 15). These components preserve low-frequency temporal dynamics and body stability. For example, $\hat{S}_{1}$ and $\hat{S}_{2}$ maintain dominant low-frequency energy at joints 2 (chest) and 3 (neck), which are critical for torso alignment and posture maintenance during the action.
Overall, this frequency-domain decomposition provides fine-grained interpretability of motion dynamics across joint hierarchies and temporal scales, confirming that the proposed spiking cross-attention fusion mechanism effectively preserves both semantic structure and dynamic variations by adaptively emphasizing functionally critical joints in the spectral domain.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) $\hat{X}$ (b) $X^{(L)}$ (c) $\hat{D}_{0}$ (d) $\hat{D}_{1}$ (e) $\hat{D}_{2}$
Refer to caption Refer to caption Refer to caption Refer to caption
(f) $\hat{S}_{0}$ (g) $\hat{S}_{1}$ (h) $\hat{S}_{2}$ (i) $\hat{S}_{3}$
Figure 10. Joint-wise frequency spectrograms of multiscale components, based on the ”drinking water” action. Each plot shows the FFT magnitude curves of all 25 joints, revealing their spectral activity at various decomposition levels.